# Stable dark matter from Pauli blocking in the degenerate fermion background with Quantum Field Theory

### v1

@@ 3.1.1 Approximate Analytic Solution
$\begin{split}\frac{d{f}_{k}}{d\tau}&\simeq q\left[2\xi_{k}\sin\left(2\tilde{\omega_{k}}\tau\right)-2\eta_{k}\cos\left(2\tilde{\omega_{k}}\tau\right)\right]\cos\tau,\\ \frac{d{\xi}_{k}}{d\tau}&\simeq q\sin(2\tilde{\omega_{k}}\tau)\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&\simeq-q\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau.\end{split}$ (3.18)
@@
$\begin{split}f_{k}(\tau)&\simeq\frac{q^{2}\left(1+12\tilde{\omega}_{k}^{2}+(-1+4\tilde{\omega}_{k}^{2})\cos(2\tau)-8\tilde{\omega}_{k}(2\tilde{\omega}_{k}\cos(\tau)\cos(2\tilde{\omega}_{k}\tau)+\sin(\tau)\sin(2\tilde{\omega}_{k}\tau))\right)}{2(1-4\tilde{\omega}_{k}^{2})^{2}},\\ \xi_{k}(\tau)&\simeq\frac{-q\left(-2\tilde{\omega}_{k}+2\tilde{\omega}_{k}\cos(\tau)\cos(2\tilde{\omega}_{k}\tau)+\sin(\tau)\sin(2\tilde{\omega}_{k}\tau)\right)}{-1+4\tilde{\omega}_{k}^{2}},\\ \eta_{k}(\tau)&\simeq\frac{q\left(\cos(2\tilde{\omega}_{k}\tau)\sin(\tau)-2\tilde{\omega}_{k}\cos(\tau)\sin(2\tilde{\omega}_{k}\tau)\right)}{-1+4\tilde{\omega}_{k}^{2}}.\end{split}$ (3.19)
@@
On the other hand, both the denominator and numerator in Eqs. (3.19) approach zero towards $\omega_{k}\sim m_{\phi}/2$. In this limit of $\omega\rightarrow m_{\phi}/2$, those asymptotic form
$\begin{split}f_{k}(\tau)&\rightarrow\frac{q^{2}}{8}(1+2\tau^{2}-\cos 2\tau+2\tau\sin 2\tau)\sim\frac{q^{2}}{4}\tau^{2},\\ \xi_{k}(\tau)&\rightarrow\frac{q}{2}(1-\cos^{2}\tau)\sim q,\\ \eta_{k}(\tau)&\rightarrow-\frac{q}{2}(\tau+\cos\tau\sin\tau)\sim-\frac{q}{2}\tau,\end{split}$ (3.20)
@@
The solution in Eq. (3.19) is invalid for $f_{k}\gtrsim 1/2$, or in other words, for the narrow resonance range
$\begin{split}\frac{m_{\phi}}{2}(1-q)\lesssim\omega_{k}\lesssim\frac{m_{\phi}}{2}(1+q).\end{split}$ (3.21)
In this regime, $f_{k}$ cannot be ignored. Instead, we find that the value of $\xi_{k}(t)$ is negligible from the numerical calculation (for example, see Figure 3). The evolution equation can be approximated as
$\begin{split}\frac{d{f}_{k}}{d\tau}&\simeq-2q\eta_{k}\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&\simeq-q\left(1-2f_{k}\right)\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau.\end{split}$ (3.22)
@@
$\begin{split}f_{k}(\tau)&\simeq\sin^{2}\left[\frac{1}{2}q\left(\tau+\cos\tau\sin\tau\right)\right]\sim\sin^{2}(q\tau/2),\\ \eta_{k}(\tau)&\simeq-\frac{1}{2}\sin\left[q(\tau+\cos\tau\sin\tau)\right]\sim-\frac{1}{2}\sin(q\tau).\end{split}$ (3.23)
These solutions match well with the numerical solutions. The solution $f_{k}(t)$ oscillates between 00 and $1$ with the period $T(\tau)=2\pi/q$ for large $\tau\gg q^{-1}$. The average over time gives $\langle f_{k}(t\rangle)=1/2$.
@@
Let’s recall the RHS of Boltzmann equation (2.14)
$\begin{split}f_{k}^{\rm coll}(t)&=\Gamma_{\psi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\delta(k-p_{\psi})\left[1-2f_{k}(t)\right].\end{split}$ (3.24)
@@
$\begin{split}\left.\langle\dot{f}_{k}\rangle\right|_{t\lesssim(qm_{\phi})^{-1}}\simeq\frac{\pi}{4}(\lambda\bar{\phi})^{2}\delta(\omega-\frac{m_{\phi}}{2})\simeq\frac{\pi^{2}n_{\phi}\Gamma_{\psi}}{p_{\psi}^{2}}\delta(k-p_{\psi}),\end{split}$ (3.25)
which is consistent with Eq. (3.24) when $f_{k}\ll 1$ This is even true for $t\gtrsim(qm_{\phi})^{-1}$ outside the resonance band.
Within the narrow band, the solution Eq. (3.23) gives a averaged value
$\begin{split}\langle f_{k}\rangle=\frac{1}{2},\end{split}$ (3.26)
which is consistent with the solution in Eq. (2.20) from Boltzmann equation without cosmic expansion for initial large $n_{\phi}$. However, the time taken for growth of $f_{\psi}$ is
$\begin{split}m_{\phi}t_{c}\sim q^{-1}\end{split}$ (3.27)
@@
$\begin{split}m_{\phi}t_{c}\simeq m_{\phi}\left(\frac{n_{\phi}(t_{i})}{p_{\psi}^{3}/(2\pi^{2})}\right)^{-1}\Gamma_{\phi}^{-1}\sim\frac{9}{4\pi^{2}}q^{-2},\end{split}$ (3.28)
@@ 3.3 With Cosmic Expansion
With taking the cosmic expansion into account, Eq. (3.10) has an additional term with Hubble expansion parameter $H$,
$\begin{split}\dot{f}_{k}&=kH\frac{\partial f_{k}}{\partial k}+i\lambda\phi(t)\left(g_{k}-g^{*}_{k}\right),\\ \dot{g}_{k}&=kH\frac{\partial g_{k}}{\partial k}-i\left[2\Omega_{k}(t)g_{k}(t)+\lambda\phi(t)\left\{\frac{K(t)^{2}}{\Omega(t)^{2}}\left(1-2f_{k}\right)+\frac{2m_{\psi}}{\Omega(t)}g_{k}\right\}\right].\end{split}$ (3.29)
With a fixed comoving momentum, $\hat{k}$, we define new functions as
$\begin{split}f_{*}(t)\equiv f_{k=K(t)}(t),\quad g_{*}(t)\equiv g_{k=K(t)}(t),\end{split}$ (3.30)
where $K(t)=\frac{a_{0}}{a(t)}\hat{k}$, and $K(t)$ is the physical momentum for the given comoving momentum $\hat{k}$ at a time $t$. Here, we choose the time of $a_{0}$ so that the magnitude of comoving momentum $\hat{k}$ becomes same as the physical momentum at that time.
@@
$\begin{split}g_{*}(t)=e^{-2i\Theta(t)}\left[\xi_{*}(t)+i\eta_{*}(t)\right],\end{split}$ (3.31)
@@
In the last equality, we used the relation in the matter-dominated Universe, $a(t)=a_{0}(t/t_{0})^{2/3}$. A time of
$K(t)=\frac{a_{0}}{a(t)}\hat{k}=\frac{m_{\phi}}{2},$ (3.33)
for a given $\hat{k}$, is expressed as
$\begin{split}\tau_{k}\equiv m_{\phi}t_{K(t)=m_{\phi}/2}=\tau_{0}\left(\frac{2\hat{k}}{m_{\phi}}\right)^{3/2}.\end{split}$ (3.34)
Now, let’s see the evolution of distribution functions for a comoving momentum $\hat{k}$ whose physical momentum is $\Omega(t_{0})=\omega_{k}$ at $t_{0}$. By using $a_{0}=a(t_{0})=1$ and $\tau\equiv m_{\phi}t$, we obtain
$\begin{split}\Omega(\tau)=\left(\frac{\tau_{0}}{\tau}\right)^{2/3}\omega_{k},\qquad\Theta(\tau)=\frac{3\omega_{k}}{m_{\phi}}\tau_{0}\left(\frac{\tau}{\tau_{0}}\right)^{1/3},\end{split}$ (3.35)
At a given time $\tau$, the period of $\sin(2\Theta(\tau))$ is about
$\begin{split}\frac{\pi}{\tilde{w}}\left(\frac{\tau_{0}}{\tau}\right)^{2/3},\end{split}$ (3.36)
which can be compared to the period $\pi/\tilde{w}_{\rm non-exp}$ in the non-expanding case with a constant $\tilde{w}_{\rm non-exp}$. Therefore, we find the correspondence of
$\begin{split}\tilde{w}_{\rm non-exp}=\tilde{w}\left(\frac{\tau_{0}}{\tau}\right)^{2/3}.\end{split}$ (3.37)
@@
$\begin{split}\phi(t)=\bar{\phi}\left(\frac{a(t)}{a_{0}}\right)^{-3/2}\cos(m_{\phi}t),\end{split}$ (3.38)
including the redshift due to the expansion of the Universe but ignoring the reduction due to decay. By using dimensionless variables,
$\begin{split}q=\frac{\lambda\bar{\phi}}{m_{\phi}},\quad\tilde{K}(t)=\frac{K(t)}{m_{\phi}},\quad\tilde{\Omega}(t)=\frac{\Omega(t)}{m_{\phi}},\end{split}$ (3.39)
and $a^{3}\propto\tau^{2}$ and taking $m_{\psi}=0$, the evolution equation of the distribution functions for comoving momentum $\hat{k}$ are simplified to be
$\begin{split}\frac{d{f}_{*}(\tau)}{d\tau}&=q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left(2\xi_{*}(\tau)\sin 2\Theta(\tau)-2\eta_{*}(\tau)\cos 2\Theta(\tau)\right)\cos\tau,\\ \frac{d{\xi}_{*}(\tau)}{d\tau}&=q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left[\left(1-2f_{*}(\tau)\right)\sin 2\Theta(\tau)\right]\cos\tau,\\ \frac{d{\eta}_{*}(\tau)}{d\tau}&=-q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left[\left(1-2f_{*}(\tau)\right)\cos 2\Theta(\tau)\right]\cos\tau,\end{split}$ (3.40)
with $\Theta(\tau)=3\tilde{\omega}_{k}\tau_{0}^{2/3}\tau^{1/3}$. Note that the equations are invariant under the scaling of $q^{\prime}=\alpha q$, $\tau_{0}^{\prime}=\alpha^{-1}\tau_{0}$ and $\omega_{k}^{\prime}=\alpha^{2/3}\omega_{k}$.
In Fig. 6, we show time evolution of the distribution $f_{*}(t)$ as well as $\xi_{*}$ and $\eta_{*}$ for a given comoving momentum $\hat{k}$ from QFT calculation in the expanding Universe. Here, we used $q=0.01$ and $\tau_{0}=10^{2}q^{-2}$. The rapid growth of $f_{*}$ happens when the corresponding physical momentum is within the narrow band in Eq. (3.21). The corresponding time for this band becomes
$\begin{split}1-\frac{2}{3}q\lesssim\frac{\tau}{\tau_{0}}\lesssim 1+\frac{2}{3}q.\end{split}$ (3.41)
@@
Note that, however, as shown in the non-expanding case of the previous section, at least $\tau\gtrsim q^{-1}$ is required for growth of $f_{k}$ even in the resonance band. Considering the time width $\Delta\tau\simeq q\tau_{0}$ in Eq. (3.41), and requiring $\Delta\tau\gtrsim q^{-1}$, at least $\tau_{0}\gg q^{-2}$ is needed for a given comoving momentum whose physics momentum becomes $m_{\phi}/2$ at $\tau_{0}$. This is again consistent with that in Boltzmann equation, Eq. (3.28). It means that the distribution function is suppressed for the comoving momentum,
$\begin{split}\hat{k}\lesssim\frac{m_{\phi}}{2}\left(\frac{1}{q^{2}\tau_{0}}\right)^{2/3},\end{split}$ (3.42)
@@
$\begin{split}k_{\rm eff}(\tau_{0})=k_{\rm eff}\left(\frac{\tau}{\tau_{0}}\right)^{2/3}\sim\sqrt{q}\left(\frac{\tau}{\tau_{0}}\right)^{1/6}m_{\phi}\ll\frac{m_{\phi}}{2},\end{split}$ (3.43)

### v2

@@ 3.1.1 Approximate Analytic Solution
$\begin{split}\frac{d{f}_{k}}{d\tau}&\simeq q\left[2\xi_{k}\sin\left(2\tilde{\omega_{k}}\tau\right)-2\eta_{k}\cos\left(2\tilde{\omega_{k}}\tau\right)\right]\cos\tau,\\ \frac{d{\xi}_{k}}{d\tau}&\simeq q\sin(2\tilde{\omega_{k}}\tau)\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&\simeq-q\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau.\end{split}$ (55)
@@
$\begin{split}f_{k}(\tau)&\simeq\frac{q^{2}\left(1+12\tilde{\omega}_{k}^{2}+(-1+4\tilde{\omega}_{k}^{2})\cos(2\tau)-8\tilde{\omega}_{k}(2\tilde{\omega}_{k}\cos(\tau)\cos(2\tilde{\omega}_{k}\tau)+\sin(\tau)\sin(2\tilde{\omega}_{k}\tau))\right)}{2(1-4\tilde{\omega}_{k}^{2})^{2}},\\ \xi_{k}(\tau)&\simeq\frac{-q\left(-2\tilde{\omega}_{k}+2\tilde{\omega}_{k}\cos(\tau)\cos(2\tilde{\omega}_{k}\tau)+\sin(\tau)\sin(2\tilde{\omega}_{k}\tau)\right)}{-1+4\tilde{\omega}_{k}^{2}},\\ \eta_{k}(\tau)&\simeq\frac{q\left(\cos(2\tilde{\omega}_{k}\tau)\sin(\tau)-2\tilde{\omega}_{k}\cos(\tau)\sin(2\tilde{\omega}_{k}\tau)\right)}{-1+4\tilde{\omega}_{k}^{2}}.\end{split}$ (56)
@@
On the other hand, both the denominator and numerator in Eqs. (56) approach zero towards $\omega_{k}\sim m_{\phi}/2$. In this limit of $\omega\rightarrow m_{\phi}/2$, those asymptotic form
$\begin{split}f_{k}(\tau)&\rightarrow\frac{q^{2}}{8}(1+2\tau^{2}-\cos 2\tau+2\tau\sin 2\tau)\sim\frac{q^{2}}{4}\tau^{2},\\ \xi_{k}(\tau)&\rightarrow\frac{q}{2}(1-\cos^{2}\tau)\sim q,\\ \eta_{k}(\tau)&\rightarrow-\frac{q}{2}(\tau+\cos\tau\sin\tau)\sim-\frac{q}{2}\tau,\end{split}$ (57)
@@
The solution in Eq. (56) is invalid for $f_{k}\gtrsim 1/2$, or in other words, for the narrow resonance range
$\begin{split}\frac{m_{\phi}}{2}(1-q)\lesssim\omega_{k}\lesssim\frac{m_{\phi}}{2}(1+q).\end{split}$ (58)
In this regime, $f_{k}$ cannot be ignored. Instead, we find that the value of $\xi_{k}(t)$ is negligible from the numerical calculation (for example, see 3). The evolution equation can be approximated as
$\begin{split}\frac{d{f}_{k}}{d\tau}&\simeq-2q\eta_{k}\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau,\\ \frac{d{\eta}_{k}}{d\tau}&\simeq-q\left(1-2f_{k}\right)\cos\left(2\tilde{\omega_{k}}\tau\right)\cos\tau.\end{split}$ (59)
@@
$\begin{split}f_{k}(\tau)&\simeq\sin^{2}\left[\frac{1}{2}q\left(\tau+\cos\tau\sin\tau\right)\right]\sim\sin^{2}(q\tau/2),\\ \eta_{k}(\tau)&\simeq-\frac{1}{2}\sin\left[q(\tau+\cos\tau\sin\tau)\right]\sim-\frac{1}{2}\sin(q\tau).\end{split}$ (60)
These solutions match well with the numerical solutions. The solution $f_{k}(t)$ oscillates between 00 and $1$ with the period $T(\tau)=2\pi/q$ for large $\tau\gg q^{-1}$. The average over time gives $\langle f_{k}(t)\rangle=1/2$.
@@
Let’s recall the RHS of Boltzmann equation (13)
$\begin{split}f_{k}^{\rm coll}(t)&=\Gamma_{\psi}n_{\phi}(t)\left(\frac{\pi^{2}}{p_{\psi}^{2}}\right)\delta(k-p_{\psi})\left[1-2f_{k}(t)\right].\end{split}$ (61)
@@
$\begin{split}\left.\langle\dot{f}_{k}\rangle\right|_{t\lesssim(qm_{\phi})^{-1}}\simeq\frac{\pi}{4}(\lambda\bar{\phi})^{2}\delta(\omega-\frac{m_{\phi}}{2})\simeq\frac{\pi^{2}n_{\phi}\Gamma_{\psi}}{p_{\psi}^{2}}\delta(k-p_{\psi}),\end{split}$ (62)
which is consistent with Eq. (61) when $f_{k}\ll 1$ This is even true for $t\gtrsim(qm_{\phi})^{-1}$ outside the resonance band.
Within the narrow band, the solution Eq. (60) gives a averaged value
$\begin{split}\langle f_{k}\rangle=\frac{1}{2},\end{split}$ (63)
which is consistent with the solution in Eq. (19) from Boltzmann equation without cosmic expansion for initial large $n_{\phi}$. However, the time taken for growth of $f_{\psi}$ is
$\begin{split}m_{\phi}t_{c}\sim q^{-1}\end{split}$ (64)
@@
$\begin{split}m_{\phi}t_{c}\simeq m_{\phi}\left(\frac{n_{\phi}(t_{i})}{p_{\psi}^{3}/(2\pi^{2})}\right)^{-1}\Gamma_{\phi}^{-1}\sim\frac{9}{4\pi^{2}}q^{-2},\end{split}$ (65)
@@ 3.3 With Cosmic Expansion
In the expanding Universe, the momentum of the fermions produced from the decay of scalar redshifts. Therefore, a specific mode which is higher than the resonance band could enter the resonance band as the Universe expands and then becomes smaller to exit the resonance band. To incorporate this redshift, we replace the time derivative in Eq. (44) as [14]
$\begin{split}\frac{d}{dt}\rightarrow\frac{\partial}{\partial t}-Hk\frac{\partial}{\partial k},\end{split}$ (66)
with the Hubble parameter $H$ and assuming the homogeneity and isotropy of the Friedmann-Robertson-Walker background. However, the RHS of Eq. (44) can be used from the flat spacetime result since the time scale of the production is much shorter than that of the Hubble expansion [15].
Taking this into account, Eq. (44) becomes,
$\begin{split}\dot{f}_{k}&=kH\frac{\partial f_{k}}{\partial k}+i\lambda\phi(t)\left(g_{k}-g^{*}_{k}\right),\\ \dot{g}_{k}&=kH\frac{\partial g_{k}}{\partial k}-i\left[2\omega_{k}g_{k}(t)+\lambda\phi(t)\left\{\frac{k^{2}}{\omega_{k}^{2}}\left(1-2f_{k}\right)+\frac{2m_{\psi}}{\omega_{k}}g_{k}\right\}\right].\end{split}$ (67)
We can simplify these equations if we use the comoving momentum as below. With a fixed comoving momentum, $\hat{k}$, we define new functions as
$\begin{split}f_{*}(t)\equiv f_{k=K(t)}(t),\quad g_{*}(t)\equiv g_{k=K(t)}(t),\end{split}$ (68)
where $K(t)=\frac{a_{0}}{a(t)}\hat{k}$ is the physical momentum for the given comoving momentum $\hat{k}$ at a time $t$. Here, we choose the time of $a_{0}$ so that the magnitude of comoving momentum $\hat{k}$ becomes the same as the physical momentum at that time.
@@
$\begin{split}g_{*}(t)=e^{-2i\Theta(t)}\left[\xi_{*}(t)+i\eta_{*}(t)\right],\end{split}$ (69)
@@
In the last equality, we used the relation in the matter-dominated Universe, $a(t)=a_{0}(t/t_{0})^{2/3}$.
@@
$\begin{split}\phi(t)=\bar{\phi}\left(\frac{a(t)}{a_{0}}\right)^{-3/2}\cos(m_{\phi}t),\end{split}$ (71)
including the redshift due to the expansion of the Universe but ignoring the reduction due to decay. Then, the evolution equations of the distribution functions Eq. (67) become for comoving momentum $\hat{k}$
$\begin{split}\frac{d{f}_{*}(\tau)}{d\tau}&=q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left(2\xi_{*}(\tau)\sin 2\Theta(\tau)-2\eta_{*}(\tau)\cos 2\Theta(\tau)\right)\cos\tau,\\ \frac{d{\xi}_{*}(\tau)}{d\tau}&=q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left[\left(1-2f_{*}(\tau)\right)\sin 2\Theta(\tau)\right]\cos\tau,\\ \frac{d{\eta}_{*}(\tau)}{d\tau}&=-q\left(\frac{\tau}{\tau_{0}}\right)^{-1}\left[\left(1-2f_{*}(\tau)\right)\cos 2\Theta(\tau)\right]\cos\tau,\end{split}$ (72)
with $q=\lambda\bar{\phi}/m_{\phi}$, and $\Theta(\tau)=3\tilde{\omega}_{k}\tau_{0}^{2/3}\tau^{1/3}$. Note that the equations are invariant under the scaling of $q^{\prime}=\alpha q$, $\tau_{0}^{\prime}=\alpha^{-1}\tau_{0}$ and $\omega_{k}^{\prime}=\alpha^{2/3}\omega_{k}$.
For a given comoving momentum $\hat{k}$, the resonance occurs when its physical momentum becomes $K(t)$ becomes the half of the scalar mass,
$K(t)=\frac{a_{0}}{a(t)}\hat{k}=\frac{m_{\phi}}{2},$ (73)
and the corresponding $\tau_{k}=m_{\phi}t_{k}$ is expressed as
$\begin{split}\tau_{k}\equiv m_{\phi}t_{K(t)=m_{\phi}/2}=\tau_{0}\left(\frac{2\hat{k}}{m_{\phi}}\right)^{3/2}.\end{split}$ (74)
Now, let’s consider a comoving momentum $\hat{k}$ whose physical momentum is $K(t_{0})=\omega_{k}$ at $t_{0}$ (or $\tau_{0}=m_{\phi}t_{0}$). By using $a_{0}=a(t_{0})=1$ and $\tau\equiv m_{\phi}t$, we obtain
$\begin{split}\Omega(\tau)=\left(\frac{\tau_{0}}{\tau}\right)^{2/3}\omega_{k},\qquad\Theta(\tau)=\frac{3\omega_{k}}{m_{\phi}}\tau_{0}\left(\frac{\tau}{\tau_{0}}\right)^{1/3}.\end{split}$ (75)
In Fig. 6, we show time evolution of the distribution $f_{*}(t)$ as well as $\xi_{*}$ and $\eta_{*}$ for a given comoving momentum $\hat{k}$ from QFT calculation in the expanding Universe. Here, we used $q=0.01$ and $\tau_{0}=10^{2}q^{-2}$. The rapid growth of $f_{*}$ happens when the corresponding physical momentum is within the narrow band in Eq. (58). The corresponding time for this band becomes
$\begin{split}1-\frac{2}{3}q\lesssim\frac{\tau}{\tau_{0}}\lesssim 1+\frac{2}{3}q.\end{split}$ (76)
@@
Note that, however, as shown in the non-expanding case of the previous section, at least $\tau\gtrsim q^{-1}$ is required for growth of $f_{k}$ even in the resonance band. Considering the time width $\Delta\tau\simeq q\tau_{0}$ in Eq. (76), and requiring $\Delta\tau\gtrsim q^{-1}$, at least $\tau_{0}\gg q^{-2}$ is needed for a given comoving momentum whose physics momentum becomes $m_{\phi}/2$ at $\tau_{0}$. This is again consistent with that in Boltzmann equation, Eq. (65). It means that the distribution function is suppressed for the comoving momentum,
$\begin{split}\hat{k}\lesssim\frac{m_{\phi}}{2}\left(\frac{1}{q^{2}\tau_{0}}\right)^{2/3},\end{split}$ (77)
@@
$\begin{split}k_{\rm eff}(\tau_{0})=k_{\rm eff}\left(\frac{\tau}{\tau_{0}}\right)^{2/3}\sim\sqrt{q}\left(\frac{\tau}{\tau_{0}}\right)^{1/6}m_{\phi}\ll\frac{m_{\phi}}{2},\end{split}$ (78)